OpenStudy (anonymous):

derivatives of (1+cos^2 7x)^3

OpenStudy (anonymous):

let f=(1+cos^2 7x)^3 =>f'=3(1+cos^2 7x)(2cos 7x)(-sin 7x)(7)

OpenStudy (anonymous):

THANKS SO MUCH! ONE OTHER QUESTION. HOW DO YOU FIND THE DERIVATIVE OF THETA? for example tan(2-theta)

OpenStudy (anonymous):

where do i start?

OpenStudy (anonymous):

let f=tan(1-theta) f'= sec^2(1-theta)(-1)

OpenStudy (anonymous):

thanks so much!!